40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 2475     Entry time: Wed Dec 11 01:29:26 2019
Author: Koji 
Type: Summary 
Category: PD QE 
Subject: Sb3513 A2P6 Dark Current / QE / Dark Noise measurement @77K 

[Raymond, Aidan, Chris, Koji]

P6 element (500um)^2

- We looked at the current amp (FEMTO) output. The amplifier saturated at the gain of 10^3 V/A. Looking at the output with a scope, we found that there is a huge 1.2MHz oscillation. Initially, we thought it is the amplifier oscillation. However, this oscillation is independent of the amplifier bandwidth when we tried the our-own made transimpedance amp.

- Shorting the cryostat chamber to the optical table made the 1.2MHz significantly reduced. Also, connecting the shield of the TEC/Laser controller made the oscillation almost invisible. This improvement allowed us to increase the amp-gain up to 10^7.

- Then the dominant RMS was 60Hz line. This was reduced by more grounding of the cable shields. The output was still dominated by the 60Hz line, but the gain could be increased to 10^8. This was sufficient for us to proceed to the careful measurements.

----

- The dark current was measured by the source meter, while the photocurrent (together with the dark current) was measured under the illumination of the ~1mW light on the PD.

- Attachment 1 shows the dependence of the dark current against the swept bias voltage. We had ~mA dark current at the room temp. So, this is ~10^5 improvement.

- Attachment 2 shows the dependence of the apparent QE against the swept bias voltage. The dark current was subtracted from the total current, to estimate the contribution of the photocurrent in the measurement.

- Attachment 3 shows the dark noise measurement at the reverse bias of ~0.6V. Up to 1kHz, the noise level was below the equivalent shotnoise level of 1mA photocurrent.

---------

All the data and python notebook in the attached zip file.

Attachment 1: Sb3513_A2P6_DarkCurrent_77K.pdf  146 kB  Uploaded Wed Dec 11 11:39:56 2019  | Hide | Hide all
Sb3513_A2P6_DarkCurrent_77K.pdf
Attachment 2: Sb3513_A2P6_QE_77K.pdf  144 kB  Uploaded Wed Dec 11 11:40:03 2019  | Hide | Hide all
Sb3513_A2P6_QE_77K.pdf
Attachment 3: Sb3513_A2P6_DarkNoise_77K.pdf  182 kB  Uploaded Wed Dec 11 11:40:09 2019  | Hide | Hide all
Sb3513_A2P6_DarkNoise_77K.pdf
Attachment 4: 191210_3513A2P6.zip  528 kB  Uploaded Wed Dec 11 11:40:17 2019
ELOG V3.1.3-